1. Using the chain rule differentiate the following function.

a. y=(2x2+4x-2)2

b. y=(x2+2x-4)2


2. Using the product rule differentiate the following functions.

a. y=2x2-x2(1/2x2+5)2

b. y=(x+2)2(x+3)2


3. Using the quotient rule differentiate the following functions.

a. y=x2/(3x+5)2

b. y=e2x/x2+2


4. Differentiate the following functions.

a. y=x2e2x

b. y=ln(x+2x+1)


5. Differentiate the following functions (with respect of to x).

a. fx=2x2+3x(4x2-2)

b. fx=ln2x2+3e2+x


6. Differentiate the following functions and find the value of x.

a. y=x2+3x(x-2)

b. y=x2-3/x+2



Thank You


1
Expert's answer
2021-09-03T12:40:09-0400

1.

a.


"y'=((2x^2+4x^{-2})^2)'"

"=2(2x^2+4x^{-2})(2x^2+4x^{-2})'"

"=4(x^2+2x^{-2})(4x-8x^{-3})"

"=\dfrac{16(x^4+2)(x^4-2)}{x^5}=\dfrac{16(x^4-4)}{x^5}"

b.


"y'=((x^2+2x-4)^2)'"

"=2(x^2+2x-4)(x^2+2x-4)'"

"=2(x^2+2x-4)(2x+2)"

"=4(x+1)(x^2+2x-4)"



2.

a.


"y'=(2x^2-x^2(\dfrac{1}{2}x^2+5)^2)'"

"=4x-2x(\dfrac{1}{2}x^2+5)^2-2x^2 (\dfrac{1}{2}x^2+5)(\dfrac{1}{2}(2x)+5)"

"=4x-2x(\dfrac{1}{2}x^2+5)^2-x^3 (x^2+10)"

"=4x-\dfrac{1}{2}x^5-10x^3-50x-x^5-10x^3"

"=-\dfrac{3x^5+40x^3+92x}{2}"

b.


"y'=((x+2)^2(x+3)^2)'"

"=2(x+2)(x+3)^2+2(x+2)^2(x+3)"

"=2(x+2)(x+3)(x+3+x+2)"

"=2(x+2)(x+3)(2x+5)"



3.

a.


"y'=(\dfrac{x^2}{(3x+5)^2})'=\dfrac{2x(3x+5)^2-2x^2(3)(3x+5)}{(3x+5)^4}"

"=\dfrac{2x(3x+5-3x)}{(3x+5)^3}=\dfrac{10x}{(3x+5)^3}"

b.


"y'=(\dfrac{e^{2x}}{x^2+2})'=\dfrac{2e^{2x}(x^2+2)-2xe^{2x}}{(x^2+2)^2}"

"=\dfrac{2e^{2x}(x^2-x+2)}{(x^2+2)^2}"



4.

a.


"y'=(x^2e^{2x})'=2xe^{2x}+2x^2e^{2x}=2x(x+1)e^{2x}"

b.


"y'=(\ln(x+2x+1))'=\dfrac{3}{3x+1}"

"y'=(\ln((x+2)(x+1))'=\dfrac{x+1+x+2}{(x+2)(x+1)}"

"=\dfrac{2x+3}{(x+2)(x+1)}"

5.

a.


"f'(x)=(2x^2+3x(4x^2-2))'"

"=(12x^3+2x^2-6x)'=36x^2+4x-6"

b.


"f'(x)=(\ln^2(x^2)+3e^{2+x})'"

"=2\ln(x^2)(\dfrac{1}{x^2})(2x)+3e^{2+x}"

"=\dfrac{8\ln|x|}{x}+3e^{2+x}"

6.

a.


"y'=(x^2+3x(x-2))'=2x+3(x-2)+3x"

"=8x-6, x\in \R"

b.


"y'=(x^2-\dfrac{3}{x+2})'=2x-\dfrac{3}{(x+2)^2}, x\not=-2"